OpenStudy (anonymous):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? A. x + 1 = 4x + 5 B. x = 4x + 5 C. x + 1 = 4x D. x + 5 = 4x + 1

OpenStudy (anonymous):

you can set both equations equal to each other, since both equations equal to y

OpenStudy (anonymous):

how do i make them equal?

OpenStudy (anonymous):

Equation A: y = x + 1 Equation B: y = 4x + 5 since they both equal to y x+1 = 4x+5

OpenStudy (anonymous):

A right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!!!